### v1

# From Federated Learning to $\mathbb{X}$-Learning: Breaking the Barriers of Decentrality Through Random Walks

Abstract
We provide our perspective on $\mathbb{X}$-Learning ($\mathbb{X}$L), a novel distributed learning architecture that generalizes and extends the concept of decentralization. Our goal is to present a vision for $\mathbb{X}$L, introducing its unexplored design considerations and degrees of freedom. To this end, we shed light on the intuitive yet non-trivial connections between $\mathbb{X}$L, graph theory, and Markov chains. We also present a series of open research directions to stimulate further research.
1 Introduction
Rapid advancement of Internet-of-Things (IoT) and widespread adoption of wireless connectivity (e.g., smart cars) have led to a surge in data-driven applications for 6G-and-beyond networks [1, 2, 3]. Collectively, modern devices generate vast amount of data, enabling the development of powerful machine learning (ML) models [4, 5]. Traditionally, these ML models are obtained using a centralized approach, requiring devices to transfer their local data to a central server. However, centralizing data can result in excessive energy consumption, high latency, and privacy concerns [6]. These challenges have fueled interest in distributed ML techniques.
1.1 Distributed ML: Federated Learning & Fog Learning
Federated learning (FedL) [7, 8, 9, 10, 11, 12, 13] is the most prominent distributed ML technique that shifts model training to data-collecting devices and uses the central servers to aggregate the locally trained models of the devices. Traditional FedL begins with a central server broadcasting a global model to devices and is characterized by two sequential steps repeated until model convergence: (i) Local training, where each device updates its model using its dataset (e.g., via stochastic gradient descent (SGD)); (ii) Global model aggregations, where the devices send their models to the central server, which aggregates the received models into a new global model (e.g., via weighted averaging) and broadcasts it to all devices. Traditional FedL fits well with the device-to-edge architecture in edge networks. However, for large-scale networks (e.g., across countries/continents), FedL must be adapted to the fog network, including nodes across the cloud-to-things continuum (e.g., the cloud, core, metro, edge, and things). This has led to the development of fog learning (FogL) [14], which addresses the extreme heterogeneity in communication and computation capabilities of fog/edge network elements for effective implementation of distributed ML.
1.2 Notion of Decentralization in FedL/FogL
The notion of decentralization has three major interpretations in distributed ML. (i) Centralized FedL [15, 16], where ML training is conducted at devices and the model aggregations are conducted at a central server through device-to-server communications (Fig. 1(A)). (ii) Semi-decentralized FogL/FedL [17, 18, 19, 20], which involves ML training at devices followed by decentralized model aggregations within local clusters of devices through one-hop device-to-device (D2D) communications and centralized aggregations across the clusters (Fig. 1(B)). (iii) Fully-decentralized FogL/FedL [21, 22, 23, 10], where ML training is conducted at devices and model aggregations are executed solely via one-hop D2D communications (Fig. 1(C)).
Figure 1: Centralized to semi/fully-decentralized FedL/FogL architectures (A)-(C), and the extension presented by the $\mathbb{X}$L (D). In centralized to semi/fully-decentralized FedL/FogL, each node has its own local ML model, whereas in $\mathbb{X}$L there are only two ML models associated with the random walkers, inducing two active sessions per training round.
1.3 From FedL/FogL to $\mathbb{X}$-Learning
In this work, we are motivated to migrate away from the above three rigid interpretations of decentralization in distributed ML, and unveil a new flexible/generic notion of decentralization through $\mathbb{X}$-Learning ($\mathbb{X}$L) that extends FedL and FogL. We thus adopted the letter $\mathbb{X}$ in its name to represent a generic family of decentralization techniques, encompassing the aforementioned notions as its special cases.
At its core, $\mathbb{X}$L is inspired by web crawlers used to index websites. The principle of $\mathbb{X}$L is to treat ML models as autonomous agents/walkers that traverse devices/nodes through D2D communications. In $\mathbb{X}$L, independent ML models use a multi-hop exploratory strategy based on random walks to traverse networks for distributed ML training.
Generalized Notion of Decentralization in $\mathbb{X}$L: $\mathbb{X}$L takes a new perspective on two core threads in FedL/FogL architectures and proposes fresh alterations to them: (i) Predetermined/Rigid Model Transfers: The interpretations of decentralization in FedL/FogL consider only predetermined/rigid one-hop device-to-server or one-hop D2D communications, where all nodes are presumed to have a unified model exchange path/trajectory. Nevertheless, $\mathbb{X}$L extends this to flexible multi-hop D2D jumps between the nodes under the presence or absence of edge servers in the network, paving the way for flexible paths/trajectories for model exchanges in complex networks. (ii) Leashed/Tied Local Model-to-Node Logic: In FedL/FogL, each device is treated as an identity with a local model (i.e., a neural network (NN)). Nevertheless, $\mathbb{X}$L shifts this identity to the ML models, treating them as autonomous agents that can jump between nodes and exploit their datasets and computation/communication resources. $\mathbb{X}$L further empowers ML models to pull each other or engage in collisions to replicate model aggregations as we will describe.
Reduction of $\mathbb{X}$L to FedL/FogL: $\mathbb{X}$L, visualized in Fig. 1(D), is a foundational distributed ML architecture that can be reduced to FedL/FogL. For example, when one walker/NN is placed at each node and set to follow a predetermined path between the node and a server, $\mathbb{X}$L reduces to FedL with centralized aggregations. Also, when one walker/NN is placed at each node and set to follow a predetermined one-hop path between the hosting node and its neighboring nodes, $\mathbb{X}$L can replicate decentralized/semi-decentralized FedL/FogL. Giving autonomy to the walkers and removing the assumption on their set/fixed paths, which in turn allows them to conduct multi-hop jumps across the nodes, opens unique opportunities for the development of flexible and generalized distributed ML methods under the umbrella of $\mathbb{X}$L.
2 Related Work, Applications, and Challenges
2.1 Related Work
$\mathbb{X}$L shares some architectural similarities to two distributed ML methods, yet possesses key distinguishing features.
(1) Gossip Learning: Gossip learning involves nodes selecting peers to exchange ML models via D2D links [24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36], where peer selection is often blind/random. In contrast, $\mathbb{X}$L features a purposeful/intelligent approach, where autonomous walkers/NNs traverse nodes over optimizable/deliberate trajectories and model training durations. While gossip learning is typically suited where blind/random D2D model transfers are sufficient (e.g., when nodes’ data is homogeneous), intelligent navigation across nodes combined with adaptable local ML model training make $\mathbb{X}$L applicable to heterogeneous/non-iid (non-independent and identically distributed) data across irregular network topologies.
(2) Random Walk-based Learning: Research has explored random-walk-based distributed learning, focusing on graph-specific tasks such as graph learning [37, 38, 39] (e.g., extracting structures/communities from a graph). Migrating from graph-specific tasks, $\mathbb{X}$L offers a fresh perspective to random walk-based learning by widening its scope to distributed ML tasks involving the training of single/multiple NNs over non-iid data scattered across a network of nodes. The closest work to $\mathbb{X}$L is a recent study that uses a single random walker in a graph, while balancing the exploration and exploitation of distributed data in FedL [40]. Compared to this, $\mathbb{X}$L is designed to be a generic distributed ML framework and introduces several degrees of freedom structured into two major innovations. (Innovation 1): Starting with a single walker (Sec. 3), we highlight the inadequacy of conventional graph theoretic centrality measures as well as Markov chain analysis for $\mathbb{X}$L, which triggers a set of novel concepts: (i) integration of node’ data and spatial qualities; (ii) elastic learning; (iii) perception of walkers; (iv) memory of walkers. (Innovation 2): We further envision multiple interacting walkers in $\mathbb{X}$L (Sec. 4), unveiling the following novel concepts: (i) optimal interactions/collaborations between walkers; (ii) attraction/repulsion between walkers. These innovations further trigger a variety of future directions that we will highlight.
2.2 Applications and Scenarios of Interest
The above innovations, combined with the generalized notion of decentrality in $\mathbb{X}$L, prove advantageous for distributed ML across a range of environments, which include:
  • Opportunistic and volatile networks: The ability of $\mathbb{X}$L walkers to jump between nodes in a D2D fashion is suitable for environments where nodes are momentarily adjacent, e.g., vehicular and aerial networks where mobile nodes have opportunistic D2D communications windows.
  • Peer-to-peer large-scale networks: $\mathbb{X}$L’s exploratory approach to learning allows for distributed ML over vast networks such as social media (e.g., Facebook), in which multiple random walkers can jump between connected/adjacent nodes (e.g., friends).
  • Resource-constrained networks: The ML training overhead within $\mathbb{X}$L does not scale with the network size, rather it scales with the number of deployed walkers. For example, with a single walker (i.e., one NN that jumps across nodes), there will only be active training at one node (i.e., one active training session at a time), regardless of the network size. This makes $\mathbb{X}$L suitable for resource-constrained networks with D2D substrates (e.g., sensor networks).
  • Autonomous networks without cellular connectivity: In remote networks or underserved/rural areas with no cellular coverage, $\mathbb{X}$L enables distributed ML, where a walker can be initialized at one of the nodes and then autonomously traverse across the nodes to learn from their data.
2.3 Challenges and Design Considerations
$\mathbb{X}$L’s novel and flexible nature, in which autonomous random walkers (i.e., NNs) are deployed as mobile learners, presents several strategic questions and challenges.
  • Which Node to Choose and How to Interact? (Spatial Challenge): In $\mathbb{X}$L, after finishing the ML model training at a node, walkers face the major challenge of selecting the next node in the network/graph to jump to. Additionally, in scenarios with multiple active walkers, they must determine ‘if’ and ‘how’ to interact with each other.
  • How Long to Stay? (Temporal Challenge): In $\mathbb{X}$L, after arriving at a node, walkers face the major challenge of how long to train at that node (e.g., how many SGD iterations should be conducted).
  • How to Integrate the Learning Result? (Learning Challenge): In $\mathbb{X}$L, after training its model on a node’s data, the walker faces a challenge of how to use this new model (e.g., whether to use it directly, replace its previous model with it, or conduct a weighted average between the two).
The Vision of $\mathbb{X}$L Finding an overarching strategy addressing the above questions, combining both network- and data-dependent considerations, is the major challenge in $\mathbb{X}$L. At first glance, it may seem that answering these questions can be informed by a mixture of concepts drawn from (i) Markov chains in non-Euclidean spaces (graph topologies), (ii) graph theory, and (iii) distributed ML. However, in the subsequent discussions, we unveil that the existing frameworks/interpretations from these fields are insufficient to fully capture the dynamics inherent to $\mathbb{X}$L. This is mainly because these frameworks are developed to focus on either the graph topology (e.g., Markov chains in non-Euclidean spaces [41, 42]) or the data quality of the nodes (e.g., majority of distributed ML literature [6]). Consequently, the efficient design and implementation of $\mathbb{X}$L is an untapped area of research, which calls for a series of follow-up works. $\star$ This perspective paper does not aim to provide a formal mathematical treatment of $\mathbb{X}$L (e.g., convergence bounds or optimization formulations), which we leave as an important direction for future research. Instead, we aim to (i) unveil the fundamental concepts and degrees of freedom in $\mathbb{X}$L, (ii) articulate them intuitively, and (iii) propose simple heuristics that can achieve notable performance gains for $\mathbb{X}$L; with the ultimate goal of triggering future research.
Henceforth, we first consider the case of $\mathbb{X}$L with a single random walker as a gateway to our discussions.
3 $\mathbb{X}$-Learning with a Single Random Walker
Considering a single walker traversing a network, the learning strategy of the walker consists of jumping/traversing between nodes and training its model on their local data using SGD. Through both conceptual reasoning and simulations, we next show that this learning strategy is tied to node importance.
Data Distribution and Network Topology All simulations use a four-layer convolutional neural network trained on CIFAR-10 [43], widely used in FedL/FogL studies. Each simulation will specify its own non-iid data partitioning. Two network topologies are considered: Caveman Graph: Used to model social/sensor networks with tight regional communities/connectivity. Random Geometric Graph (RGG): Used to model D2D links over wireless networks [44]. $\star$ These graphs were intentionally chosen to examine $\mathbb{X}$L in network topologies with highly unbalanced (caveman) and semi-unbalanced (RGG) connectivity.
3.1 Node Importance/Centrality
In both $\mathbb{X}$L and FedL/FogL, the quality of the local data at a node is a key to determining its importance: the more representative the local data is of the global dataset (i.e., the union of the nodes’ local data), the higher its quality. Although this metric alone may suffice for FedL/FogL, it is insufficient for $\mathbb{X}$L. To demonstrate this, consider a node with poor data quality serving as a connective bridge or hub between multiple cliques of nodes, each containing high-quality data. Here, while the node itself may not offer high-quality data for training, its network centrality (i.e., spatial quality) can make it highly valuable for the walker to visit as a gateway for reaching nodes with higher-quality data. As a result, in $\mathbb{X}$L, if we aim to model the jumps/traversal between the nodes for a walker through a probability transition matrix (similar to Markov chains over graphs), this matrix should be a function of both the nodes’ data qualities and their network centralities. To achieve this goal, the concept of centrality in graph/network theory (e.g., measured via degree, betweenness, eigenvalue, and closeness metrics [45, 41, 42]) should be integrated to the concept of data quality (e.g., measured via the disparity of gradients generated on the local data of a node versus the global dataset [6]). By jointly considering spatial and data quality metrics, walkers in $\mathbb{X}$L environments follow trajectories that are adaptive to nodes’ data distribution and network topology. However, selecting a suitable realization for each metric and determining how to combine them (e.g., weighted sum of data quality and spatial quality) remains a non-trivial and open research challenge.
Figure 2: Performance of random walkers under different node traversal strategies. Referring to Sec. 3.1, among methods with fixed/static node importance values (red, blue, green, magenta, and yellow curves using static sampling or fixed $\alpha\in{0,0.5,1}$), our approach with non-binary weighting of data and spatial quality metrics (yellow curve, $\alpha=0.5$) yields the best performance. Furthermore, referring to Sec. 3.3, our method with dynamic node importance (black curve using a time-varying $\alpha^{\mathsf{Inst}}$ based on current model accuracy) outperforms all the static strategies.
To validate the above-described approach to node importance, we conduct a simulation and present it in Fig. 2. Here, we consider a caveman graph comprising $50$ nodes clustered into $8$ cliques. The nodes’ data distribution is non-iid with $98\%$ of nodes only containing data samples from $1$ or $2$ of the $10$ labels in CIFAR-10 dataset. We assume that the walker traverses the nodes according to various strategies, including uniform random sampling (blue curve) and Metropolis-Hastings sampling [46] (green curve) — a well-known technique in Markov chains over graphs. We consider a simple realization of our node importance approach, where the importance of node $i$ is computed as a weighted sum: $\widetilde{\mathfrak{c}}_{i}=\alpha\cdot d_{i}\cdot\ell_{i}+(1-\alpha)\cdot c_{i}$. Here, the weighting parameter $\alpha\in[0,1]$ controls the walker’s preference between data and spatial quality. In this formulation, $d_{i}$ and $\ell_{i}$ represent the fractions ($d_{i},\ell_{i}\in[0,1]$) of global dataset and labels present at node $i$ (i.e., capturing data quality), respectively, and $c_{i}$ is a conventional measure of the node’s centrality (chosen to be betweenness centrality [41, 42], capturing the spatial quality). At each node, the transition probability of the walker to the neighboring nodes is defined proportional to their importance values ($\{\widetilde{\mathfrak{c}}_{i}\}$). Our scheme (yellow curve) that obtains node importance as a balanced average ($\alpha=0.5$) between data and spatial quality, significantly outperforms other strategies, including the ones that only consider spatial quality ($\alpha=0$) and data quality ($\alpha=1$). This balanced value of $\alpha$ may not hold across all environments, but the results highlight the significance of considering network topology alongside the nodes’ data quality. We later (Sec. 3.3) show that we can further obtain dynamic transition matrices to achieve a better performance (i.e., the black curve in Fig. 2).
Motivated by the above results, we conjecture that optimal node traversal in $\mathbb{X}$L demands moving beyond conventional graph exploration and distributed ML approaches, such as Markov chain-based methods that focus solely on network topology while ignoring node data quality, and FedL/FogL techniques that emphasize only data quality while overlooking dynamic multi-hop graph exploration via random walks. A formal treatment of this conjecture is left for future work.
3.2 Elastic Learning
In $\mathbb{X}$L, the walker trains on the data of the node it visits. Seemingly counter-intuitive, remaining at every node for extensive training and model refinement (i.e., a greater number of SGD iterations) is not always optimal. Instead, the decision on the model training duration should depend on both the quality of the node’s data and the current state of the walker’s model. For instance, if the walker’s model is already well-trained, performing many updates/SGDs on a node with low-quality (e.g., biased) data can degrade its performance, while more updates/SGDs can be conducted on nodes with high-quality data. The need for such dynamic range of update/SGD durations across nodes reveals a key degree of freedom in $\mathbb{X}$L, which we call elasticity of local training.
In Fig. 3, we show the elasticity of local training through a simulation over an RGG with $100$ nodes, $90\%$ of which have only $1$ or $2$ labels of CIFAR-10 dataset. In the three baselines, the walker is assumed to be inelastic and performing a fixed number of SGD iterations at each node ($20$, $40$, and $60$ iterations). Also, we consider an elastic walker (red curve), where its number of SGD iterations $x_{i}$ at each node $i$ is tuned according to the node’s data quality, up to a maximum of $x^{\textrm{Max}}=20$ using a sigmoid-like function $x_{i}=\frac{x^{\textrm{Max}}}{1+\exp({-\tau_{1}\times Q_{i}})}$, where $\tau_{1}=10$ is a temperature parameter, and $Q_{i}$ is a relative measure of the node $i$’s data quality defined by $Q_{i}=\ell_{i}\times{d}_{i}^{\tau_{2}\times(1-{d_{i}})}$ with $\tau_{2}=0.4$ and the rest of the variables ($d_{i}$$\ell_{i}$) are the same as Sec. 3.1. The form of this function allows for a sharp reduction in the number of SGD iterations executed at nodes with low-quality data, while its hyperparameters were chosen experimentally. The results unveil that the elastic walker performs markedly better than the inelastic ones, all while training on significantly fewer data samples. Other formulations for tuning SGD iterations according to node data quality are open for exploration.
Figure 3: Performance comparison between our elastic random walker, where the number of SGD iterations is adaptively scaled based on node data quality, and methods with fixed SGD iteration numbers. Our elastic walker (red curves in the top and bottom plots) outperforms all baselines with static SGD iteration numbers (top plot) despite performing fewer SGD updates (bottom plot).
3.3 Perception-Aware Random Walk Design
We next introduce random walk perception resolution in $\mathbb{X}$L, where the walker adjusts its node traversal based on its model accuracy, shifting from static node importance-based traversals (Sec. 3.1) to accuracy-aware dynamic exploration. Specifically, early in training, with low accuracy, the walker has low perception resolution and treats most nodes as important for broad exploration. As its accuracy improves, the walker’s resolution sharpens, making it more selective, favoring nodes with high-quality data and avoiding low-quality ones to prevent performance degradation.
This further unveils a perception-aware trajectory design, where the walker’s trajectory changes over time, suggesting a departure from time-homogeneous to time-inhomogeneous Markov Chains over graphs for $\mathbb{X}$L. To show this, we redefine the node importance as a function of the walker’s model accuracy at jump/time $t$ as $\widetilde{\mathfrak{c}}_{i}(t)=\alpha^{\mathsf{Inst}}(t)\cdot d_{i}\cdot\ell_{i}+(1-\alpha^{\mathsf{Inst}}(t))\cdot c_{i}$, where compared to the fixed value of $\alpha$ used in Sec. 3.1, here $\alpha^{\mathsf{Inst}}(t)\in[\alpha^{\mathsf{Min}}{=}0.10\leavevmode\nobreak\ ,\leavevmode\nobreak\ \alpha^{\mathsf{Max}}{=}0.85]$ is a dynamic/instantaneous weighting coefficient. We model $\alpha^{\mathsf{Inst}}(t)$ as a simple function of the walker’s instantaneous accuracy $\text{A}^{\mathsf{Inst}}(t)$ as
$\alpha^{\mathsf{Inst}}(t)=\alpha^{\mathsf{Min}}+\big{(}(\text{A}^{\mathsf{Inst}}(t)-\text{A}^{\mathsf{Min}})\cdot(\alpha^{\mathsf{Max}}-\alpha^{\mathsf{Min}})\big{)}\big{/}\big{(}\text{A}^{\mathsf{Max}}-\text{A}^{\mathsf{Min}}\big{)}$, where $\text{A}^{\mathsf{Min}}=0.1$ and $\text{A}^{\mathsf{Max}}=0.8$ capture the estimated minimum and maximum achievable accuracy of the walker. Subsequently, at each node, the transition probability of the walker to the neighboring nodes is defined proportional to their time-varying importance values ($\{\widetilde{\mathfrak{c}}_{i}(t)\}$), resulting in a time-inhomogeneous Markov Chain over the network graph. Inspecting Fig. 2, the above-described perception-aware walker (black curve) exhibits a superior performance. We leave exploration of more advanced formulations of perception-aware trajectory design as future work.
Figure 4: The performance comparisons between a memory-enabled random walker and a walker with no memory, both following the same node traversal strategy. Initially, memory is disabled (red region) to accelerate early training. As training progresses, memory influence gradually increases (green and blue regions), helping to mitigate model bias and yielding notable performance gains; especially in the middle to late stages of model training.
3.4 Random Walks with Flexible Memory
A key challenge for a random walker is forgetting patterns learned from previously visited nodes as it continues updating its model during new node visits, mirroring the catastrophic forgetting phenomenon in ML. To address this, we propose flexible-memory walkers that maintain two model instances: a stale model (SM) and an instantaneous model (IM). Initially, both models are synchronized. Upon visiting a node, only the IM is updated using its local data. Afterwards, the walker averages the SM and IM to produce an aggregated model and uses it to synchronize both models. This process helps the IM retain knowledge from earlier visits.
To verify this strategy, we conduct a simulation on an RGG with $500$ nodes, $90\%$ of which possessing $1$ or $2$ of the CIFAR-10 labels. The walker is presumed to have an IM, $\bm{w}^{\mathsf{IM}}$, and an SM, $\bm{w}^{\mathsf{SM}}$. After arriving at each node, $\bm{w}^{\mathsf{IM}}$ is first updated via conducting $5$ SGDs on the node’s data. Then, the IM is updated via aggregation with the SM as $\bm{w}^{\mathsf{IM}}\leftarrow(1-\beta)\bm{w}^{\mathsf{IM}}+\beta\bm{w}^{\mathsf{SM}}$; the updated IM is further used to refine the SM as $\bm{w}^{\mathsf{SM}}\leftarrow\bm{w}^{\mathsf{IM}}$. Here, $\beta\in[0,1]$ captures the importance of the SM. The performance improvements for the memory-enabled walker are notable in Fig. 4, where in the early stages (red area), memory was disabled to accelerate the initial learning and its effect was gradually increased in stages (green and purple areas). For demonstrations, the value of $\beta$ at each jump was obtained through a line search across the interval written in each shaded region in Fig. 4. A more generalized formulation for memory-enabled walker is left for future work.
Future Research Directions The exploration of $\mathbb{X}$L with a single walker opens up various future directions, discussed below. Connection to Graph Signal Processing (GSP): We suggest exploring node importance metrics for $\mathbb{X}$L inspired by core signal processing methods, such as GSP [47, 48, 49, 50, 51]. GSP incorporates graph topology into node signal representation through eigen-decomposition of the graph Laplacian. Developing a unified metric for node importance in $\mathbb{X}$L via GSP calls for (i) defining appropriate signals for the nodes based on their data, (ii) studying interpretations of eigen-decomposition for $\mathbb{X}$L, and (iii) leveraging GSP-inspired learning mechanisms (e.g., graph neural networks) for $\mathbb{X}$L. ML Convergence under Time-Varying Transition Matrices: Characterizing the ML convergence under time-varying transition matrices is crucial and will involve obtaining ML convergence bounds for convex and non-convex loss functions, considering the impacts of random walk trajectory on its model convergence. Comparison with Traditional Markov Chain Designs: Analyzing how the dynamic design of transition matrices differs from time-homogeneous Markov chains considering traditional metrics (e.g., covering time and mixing time [52, 53]) is interesting. This involves demonstrating ‘how’ and ‘why’ traditional random walk visiting patterns may lead to inferior performance when compared to $\mathbb{X}$L-specific strategies. Resource-Aware $\mathbb{X}$L: To derive the optimal walker trajectory and training durations at nodes, future work can formulate optimization problems with competing objective functions, comprising the walker’s ML performance, communication/traversal latency, and energy consumption, under realistic transition costs (e.g., energy and delay) between nodes.
Figure 5: Impact of the number of random walkers on model convergence. The top plot shows the instantaneous average performance across all walkers, where increasing the number of walkers initially boosts performance; however, beyond a certain point (e.g., from 10 to 14 walkers), the improvements become marginal. The bottom plot presents the walkers’ final accuracies, highlighting an initial sharp gain followed by diminishing gains as more walkers are deployed.
4 $\mathbb{X}$-Learning with Multiple Random Walkers
Expanding $\mathbb{X}$L, we consider scenarios involving multiple random walkers, opening up new research opportunities. Here, the dynamics of learning and the interaction between walkers add new layers of complexity. For instance, multiple walkers can have overlapping traversal patterns and potential collaborative or competitive dynamics as detailed next.
4.1 Collaboration Among Walkers
Having multiple random walkers independently exploring the network can be inefficient, motivating us to study interactions between the walkers. Drawing from physics concepts on interacting particles [54, 55, 56, 57, 58] and epidemic spreading [59, 60, 61, 62, 63, 64], we explore cooperation between random walkers through information transfer. Specifically, when two walkers meet at a node, they “collide” and aggregate their models (e.g., via weighted averaging their models based on the number of seen datapoints). To show this, we simulate $\mathbb{X}$L over a caveman graph with $10$ cliques, each containing $100$ nodes, where each clique is dominated by the presence of a single label of CIFAR-10. All walkers select their next nodes to jump uniformly at random; however, every $10$ jumps, each pair of walkers aggregate by visiting a predetermined node. Fig. 5 shows that collisions and information exchange are beneficial. Further, it shows the diminishing reward of the number of deployed walkers. Investigating the factors influencing this diminishing reward is an exciting research direction.
Figure 6: Impact of inter-walker attraction on model performance. As the attraction between walkers increases (i.e., higher values of AA, leading to shorter inter-collision intervals), the walkers interact more frequently, which results in mitigating model bias and an improved performance. For comparison, the red curve illustrates the (hypothetical) upper bound achieved by centrally aggregating the walkers’ models after every jump.
4.2 Attraction/Repulsion Between Multiple Random Walkers
We next aim to control the frequency of collisions between walkers. Consider a scenario in which two walkers start at the same node and follow divergent paths. These walkers may develop distinct (biased) models as they traverse and train across different nodes. When walkers eventually collide, the aggregated model can potentially exhibit a low performance if their models are highly biased/separated. Drawing inspiration from charged particles, we propose mechanisms to periodically pull/attract walkers, enabling more frequent aggregations and reducing bias. Essentially, the longer two walkers travel without colliding, the stronger their polarity and thus attraction, pulling them towards each other for eventual aggregation. After a collision, walkers briefly lose attraction, preventing redundant aggregations.
To demonstrate, we set up a caveman graph comprising $64$ nodes grouped into $8$ cliques. In this scenario, $8$ random walkers are deployed, one in each clique and their transition matrices are tuned such that they solely explore their own cliques. The level of attraction between the walkers was dictated by a configurable parameter $A$, higher values of which lead to more frequent collisions (i.e., shorter inter-collision intervals). In particular, we defined the attraction between each pair of walkers $r$ and $r^{\prime}$ at jump instant $t$ via the probability $p_{r\leftrightarrow r^{\prime}}{(t)}=\min\{1,\exp{(A\cdot\mathcal{T}_{r,r^{\prime}}{(t)})}\}$, where $\mathcal{T}_{r,r^{\prime}}(t)$ represents the elapsed time since the walker pair collided. For each walker $r$, its transition matrix is extended such that attractions between $r$ and all other walkers $r^{\prime}$ are included, i.e., $p_{r\leftrightarrow r^{\prime}}(t),\forall r^{\prime}$. The results are presented in Fig. 6: as parameter $A$ increases and collisions become more frequent, the walkers’ performance improves, approaching a hypothetical best-case scenario with centralized aggregation of walkers after every jump (labeled as ‘uplink’).
Future Research Directions The above exploration of $\mathbb{X}$L with multiple random walkers opens up a series of future directions: Incorporating Caching: The need for physical collisions between walkers can be reduced by assuming that walkers leave a trace/footprint of their ML models on nodes. This trace gradually fades over time to signify the staleness of the left ML model. If a new walker arrives at a node where a trace is still fresh/reliable, it can aggregate its ML model with this trace, while accounting for the age of the trace. From Single-Task to Multi-Task Learning: The walkers can potentially train different ML tasks, introducing multi-task learning [65, 66, 67, 68, 69] into $\mathbb{X}$L. This opens up exploration of (i) aggregation between the walkers based on their task similarity, (ii) optimal walker-to-task allocation, and (iii) resource optimization (e.g., number of SGD iterations of walkers). Security of Random Walkers and their Byzantine Resilience: We suggest studying how exploiting the parameters of random walkers during node traversal (e.g., via implementing an inference attack) in different network neighborhoods (e.g., areas with high vs. low data quality) can reveal information about the data of their previously visited nodes. Additionally, exploring the walkers’ robustness against Byzantine attackers (e.g., malicious random walkers that intentionally aggregate with legitimate walkers to degrade their performance) is an interesting direction. Redefining the Notion of Centrality and Perception: Incorporating multiple random walkers necessitates reevaluating node centrality and walkers’ perceptions defined in Sec. 3. These revamps should be done to avoid redundancy in exploration by assigning different centrality values to a node across walkers.
5 Conclusion
We introduced $\mathbb{X}$L, a novel framework that expands distributed ML by treating ML models as autonomous agents capable of traversing the network through flexible, multi-hop strategies. The presented formulations and empirical results demonstrated the significant potential of $\mathbb{X}$L. A series of future research directions were discussed, broadening the horizon of $\mathbb{X}$L.
Acknowledgements
Not applicable
Funding
S. Hosseinaliour was partially funded by NSF 2421761 while working on this paper.
Competing interests
The authors declare no competing interests.
Additional information
Correspondence and requests for materials should be addressed to S. Hosseinalipour.

### v2

@@
 Abstract
-We provide our perspective on $\mathbb{X}$-Learning ($\mathbb{X}$L), a novel distributed learning architecture that generalizes and extends the concept of decentralization.
-Our goal is to present a vision for $\mathbb{X}$L, introducing its unexplored design considerations and degrees of freedom.
-To this end, we shed light on the intuitive yet non-trivial connections between $\mathbb{X}$L, graph theory, and Markov chains.
-We also present a series of open research directions to stimulate further research.
-1 Introduction
-Rapid advancement of Internet-of-Things (IoT) and widespread adoption of wireless connectivity (e.g., smart cars) have led to a surge in data-driven applications for 6G-and-beyond networks [1, 2, 3].
-Collectively, modern devices generate vast amount of data, enabling the development of powerful machine learning (ML) models [4, 5].
+Modern wireless networks have witnessed an unprecedented surge in data generation across distributed devices.
+This data can be used to empower a variety of downstream applications (e.g., autonomous driving) through machine learning (ML) techniques.
+Nevertheless, traditional centralized ML techniques face significant challenges in wireless networks, as they require pooling data from wireless devices at a central server, leading to high energy consumption, latency, and privacy risks.
+To address these issues, distributed ML techniques, such as federated learning (FedL) and fog learning (FogL), have emerged.
+Although these techniques enable model training directly on devices, they are often rigid in their decentralization strategies, reducing their adaptability to a wide range of wireless networks (e.g., opportunistic and volatile networks).
+This paper introduces $\mathbb{X}$-Learning ($\mathbb{X}$L), a novel distributed ML architecture that generalizes and extends the concept of decentralization.
+Inspired by web crawlers, $\mathbb{X}$L treats ML models as autonomous agents that navigate the network using a multi-hop random walk-based strategy, learning from diverse data across nodes.
+This flexible framework can adapt to various network conditions and topologies.
+Our goal is to present a vision for $\mathbb{X}$L, introducing its nuanced design considerations and unexplored degrees of freedom.
+To this end, we explore and unveil the foundations of $\mathbb{X}$L in both single random walker and multiple random walker settings.
+We also propose a set of new notions and concepts that yield novel insights on the working mechanism of $\mathbb{X}$L.
+We further shed light on the intuitive yet non-trivial connections between $\mathbb{X}$L, graph theory, and Markov chains.
+Finally, we present a series of open research directions that can trigger a variety of follow-up works.
+I Introduction
+Rapid advancement of Internet-of-Things (IoT) and widespread adoption of wireless connectivity (e.g., smart cars) have led to a surge in data-driven applications for 6G-and-beyond networks.
+Collectively, modern devices generate vast amount of data, enabling the development of powerful machine learning (ML) models.
 Traditionally, these ML models are obtained using a centralized approach, requiring devices to transfer their local data to a central server.
@@
 These challenges have fueled interest in distributed ML techniques.
-1.1 Distributed ML: Federated Learning & Fog Learning
-Federated learning (FedL) [7, 8, 9, 10, 11, 12, 13] is the most prominent distributed ML technique that shifts model training to data-collecting devices and uses the central servers to aggregate the locally trained models of the devices.
+I-A Distributed ML: Federated Learning & Fog Learning
+Federated learning (FedL) [2] is the most prominent distributed ML technique that shifts model training to data-collecting devices and uses the central servers to aggregate the locally trained models of the devices.
 Traditional FedL begins with a central server broadcasting a global model to devices and is characterized by two sequential steps repeated until model convergence: (i) Local training, where each device updates its model using its dataset (e.g., via stochastic gradient descent (SGD)); (ii) Global model aggregations, where the devices send their models to the central server, which aggregates the received models into a new global model (e.g., via weighted averaging) and broadcasts it to all devices.
@@
 This has led to the development of fog learning (FogL) [3], which addresses the extreme heterogeneity in communication and computation capabilities of fog/edge network elements for effective implementation of distributed ML.
-1.2 Notion of Decentralization in FedL/FogL
+I-B Notion of Decentralization in FedL/FogL
 The notion of decentralization has three major interpretations in distributed ML.
-(i) Centralized FedL [15, 16], where ML training is conducted at devices and the model aggregations are conducted at a central server through device-to-server communications (Fig.
+(i) Centralized FedL, where ML training is conducted at devices and the model aggregations are conducted at a central server through device-to-server communications (Fig.
 1(A)).
-(ii) Semi-decentralized FogL/FedL [17, 18, 19, 20], which involves ML training at devices followed by decentralized model aggregations within local clusters of devices through one-hop device-to-device (D2D) communications and centralized aggregations across the clusters (Fig.
+(ii) Semi-decentralized FogL/FedL, which involves ML training at devices followed by decentralized model aggregations within local clusters of devices through one-hop device-to-device (D2D) communications and centralized aggregations across the clusters [4] (Fig.
 1(B)).
-(iii) Fully-decentralized FogL/FedL [21, 22, 23, 10], where ML training is conducted at devices and model aggregations are executed solely via one-hop D2D communications (Fig.
+(iii) Fully-decentralized FogL/FedL, where ML training is conducted at devices and model aggregations are executed solely via one-hop D2D communications [5, 2] (Fig.
 1(C)).
@@
 In centralized to semi/fully-decentralized FedL/FogL, each node has its own local ML model, whereas in $\mathbb{X}$L there are only two ML models associated with the random walkers, inducing two active sessions per training round.
-1.3 From FedL/FogL to $\mathbb{X}$-Learning
+I-C From FedL/FogL to $\mathbb{X}$-Learning
 In this work, we are motivated to migrate away from the above three rigid interpretations of decentralization in distributed ML, and unveil a new flexible/generic notion of decentralization through $\mathbb{X}$-Learning ($\mathbb{X}$L) that extends FedL and FogL.
@@
 Giving autonomy to the walkers and removing the assumption on their set/fixed paths, which in turn allows them to conduct multi-hop jumps across the nodes, opens unique opportunities for the development of flexible and generalized distributed ML methods under the umbrella of $\mathbb{X}$L.
-2 Related Work, Applications, and Challenges
-2.1 Related Work
+II Related Work, Applications, and Challenges
+II-A Related Work
 $\mathbb{X}$L shares some architectural similarities to two distributed ML methods, yet possesses key distinguishing features.
-(1) Gossip Learning: Gossip learning involves nodes selecting peers to exchange ML models via D2D links [24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36], where peer selection is often blind/random.
+(1) Gossip Learning: Gossip learning involves nodes selecting peers to exchange ML models via D2D links [6], where peer selection is often blind/random.
 In contrast, $\mathbb{X}$L features a purposeful/intelligent approach, where autonomous walkers/NNs traverse nodes over optimizable/deliberate trajectories and model training durations.
 While gossip learning is typically suited where blind/random D2D model transfers are sufficient (e.g., when nodes’ data is homogeneous), intelligent navigation across nodes combined with adaptable local ML model training make $\mathbb{X}$L applicable to heterogeneous/non-iid (non-independent and identically distributed) data across irregular network topologies.
-(2) Random Walk-based Learning: Research has explored random-walk-based distributed learning, focusing on graph-specific tasks such as graph learning [37, 38, 39] (e.g., extracting structures/communities from a graph).
+(2) Random Walk-based Learning: Research has explored random-walk-based distributed learning, focusing on graph-specific tasks such as graph learning [7] (e.g., extracting structures/communities from a graph).
 Migrating from graph-specific tasks, $\mathbb{X}$L offers a fresh perspective to random walk-based learning by widening its scope to distributed ML tasks involving the training of single/multiple NNs over non-iid data scattered across a network of nodes.
@@
 (Innovation 1): Starting with a single walker (Sec.
-3), we highlight the inadequacy of conventional graph theoretic centrality measures as well as Markov chain analysis for $\mathbb{X}$L, which triggers a set of novel concepts: (i) integration of node’ data and spatial qualities; (ii) elastic learning; (iii) perception of walkers; (iv) memory of walkers.
+III), we highlight the inadequacy of conventional graph theoretic centrality measures as well as Markov chain analysis for $\mathbb{X}$L, which triggers a set of novel concepts: (i) integration of node’ data and spatial qualities; (ii) elastic learning; (iii) perception of walkers; (iv) memory of walkers.
 (Innovation 2): We further envision multiple interacting walkers in $\mathbb{X}$L (Sec.
-4), unveiling the following novel concepts: (i) optimal interactions/collaborations between walkers; (ii) attraction/repulsion between walkers.
+IV), unveiling the following novel concepts: (i) optimal interactions/collaborations between walkers; (ii) attraction/repulsion between walkers.
 These innovations further trigger a variety of future directions that we will highlight.
-2.2 Applications and Scenarios of Interest
+II-B Applications and Scenarios of Interest
 The above innovations, combined with the generalized notion of decentrality in $\mathbb{X}$L, prove advantageous for distributed ML across a range of environments, which include:
@@
   • Autonomous networks without cellular connectivity: In remote networks or underserved/rural areas with no cellular coverage, $\mathbb{X}$L enables distributed ML, where a walker can be initialized at one of the nodes and then autonomously traverse across the nodes to learn from their data.
-2.3 Challenges and Design Considerations
+II-C Challenges and Design Considerations
 $\mathbb{X}$L’s novel and flexible nature, in which autonomous random walkers (i.e., NNs) are deployed as mobile learners, presents several strategic questions and challenges.
@@
 Consequently, the efficient design and implementation of $\mathbb{X}$L is an untapped area of research, which calls for a series of follow-up works.
-$\star$ This perspective paper does not aim to provide a formal mathematical treatment of $\mathbb{X}$L (e.g., convergence bounds or optimization formulations), which we leave as an important direction for future research.
+$\star$ This magazine paper does not aim to provide a formal mathematical treatment of $\mathbb{X}$L (e.g., convergence bounds or optimization formulations), which we leave as an important direction for future research.
 Instead, we aim to (i) unveil the fundamental concepts and degrees of freedom in $\mathbb{X}$L, (ii) articulate them intuitively, and (iii) propose simple heuristics that can achieve notable performance gains for $\mathbb{X}$L; with the ultimate goal of triggering future research.
 Henceforth, we first consider the case of $\mathbb{X}$L with a single random walker as a gateway to our discussions.
-3 $\mathbb{X}$-Learning with a Single Random Walker
+III $\mathbb{X}$-Learning with a Single Random Walker
 Considering a single walker traversing a network, the learning strategy of the walker consists of jumping/traversing between nodes and training its model on their local data using SGD.
@@
 $\star$ These graphs were intentionally chosen to examine $\mathbb{X}$L in network topologies with highly unbalanced (caveman) and semi-unbalanced (RGG) connectivity.
-3.1 Node Importance/Centrality
+III-A Node Importance/Centrality
 In both $\mathbb{X}$L and FedL/FogL, the quality of the local data at a node is a key to determining its importance: the more representative the local data is of the global dataset (i.e., the union of the nodes’ local data), the higher its quality.
@@
 As a result, in $\mathbb{X}$L, if we aim to model the jumps/traversal between the nodes for a walker through a probability transition matrix (similar to Markov chains over graphs), this matrix should be a function of both the nodes’ data qualities and their network centralities.
-To achieve this goal, the concept of centrality in graph/network theory (e.g., measured via degree, betweenness, eigenvalue, and closeness metrics [45, 41, 42]) should be integrated to the concept of data quality (e.g., measured via the disparity of gradients generated on the local data of a node versus the global dataset [6]).
+To achieve this goal, the concept of centrality in graph/network theory (e.g., measured via degree, betweenness, eigenvalue, and closeness metrics [9, 10]) should be integrated to the concept of data quality (e.g., measured via the disparity of gradients generated on the local data of a node versus the global dataset [1]).
 By jointly considering spatial and data quality metrics, walkers in $\mathbb{X}$L environments follow trajectories that are adaptive to nodes’ data distribution and network topology.
@@
 Referring to Sec.
-3.1, among methods with fixed/static node importance values (red, blue, green, magenta, and yellow curves using static sampling or fixed $\alpha\in{0,0.5,1}$), our approach with non-binary weighting of data and spatial quality metrics (yellow curve, $\alpha=0.5$) yields the best performance.
+III-A, among methods with fixed/static node importance values (red, blue, green, magenta, and yellow curves using static sampling or fixed $\alpha\in{0,0.5,1}$), our approach with non-binary weighting of data and spatial quality metrics (yellow curve, $\alpha=0.5$) yields the best performance.
 Furthermore, referring to Sec.
-3.3, our method with dynamic node importance (black curve using a time-varying $\alpha^{\mathsf{Inst}}$ based on current model accuracy) outperforms all the static strategies.
+III-C, our method with dynamic node importance (black curve using a time-varying $\alpha^{\mathsf{Inst}}$ based on current model accuracy) outperforms all the static strategies.
 To validate the above-described approach to node importance, we conduct a simulation and present it in Fig.
@@
 We later (Sec.
-3.3) show that we can further obtain dynamic transition matrices to achieve a better performance (i.e., the black curve in Fig.
+III-C) show that we can further obtain dynamic transition matrices to achieve a better performance (i.e., the black curve in Fig.
 2).
@@
 A formal treatment of this conjecture is left for future work.
-3.2 Elastic Learning
+III-B Elastic Learning
 In $\mathbb{X}$L, the walker trains on the data of the node it visits.
@@
 Also, we consider an elastic walker (red curve), where its number of SGD iterations $x_{i}$ at each node $i$ is tuned according to the node’s data quality, up to a maximum of $x^{\textrm{Max}}=20$ using a sigmoid-like function $x_{i}=\frac{x^{\textrm{Max}}}{1+\exp({-\tau_{1}\times Q_{i}})}$, where $\tau_{1}=10$ is a temperature parameter, and $Q_{i}$ is a relative measure of the node $i$’s data quality defined by $Q_{i}=\ell_{i}\times{d}_{i}^{\tau_{2}\times(1-{d_{i}})}$ with $\tau_{2}=0.4$ and the rest of the variables ($d_{i}$$\ell_{i}$) are the same as Sec.
-3.1.
+III-A.
 The form of this function allows for a sharp reduction in the number of SGD iterations executed at nodes with low-quality data, while its hyperparameters were chosen experimentally.
@@
 Our elastic walker (red curves in the top and bottom plots) outperforms all baselines with static SGD iteration numbers (top plot) despite performing fewer SGD updates (bottom plot).
-3.3 Perception-Aware Random Walk Design
+III-C Perception-Aware Random Walk Design
 We next introduce random walk perception resolution in $\mathbb{X}$L, where the walker adjusts its node traversal based on its model accuracy, shifting from static node importance-based traversals (Sec.
-3.1) to accuracy-aware dynamic exploration.
+III-A) to accuracy-aware dynamic exploration.
 Specifically, early in training, with low accuracy, the walker has low perception resolution and treats most nodes as important for broad exploration.
@@
 To show this, we redefine the node importance as a function of the walker’s model accuracy at jump/time $t$ as $\widetilde{\mathfrak{c}}_{i}(t)=\alpha^{\mathsf{Inst}}(t)\cdot d_{i}\cdot\ell_{i}+(1-\alpha^{\mathsf{Inst}}(t))\cdot c_{i}$, where compared to the fixed value of $\alpha$ used in Sec.
-3.1, here $\alpha^{\mathsf{Inst}}(t)\in[\alpha^{\mathsf{Min}}{=}0.10\leavevmode\nobreak\ ,\leavevmode\nobreak\ \alpha^{\mathsf{Max}}{=}0.85]$ is a dynamic/instantaneous weighting coefficient.
+III-A, here $\alpha^{\mathsf{Inst}}(t)\in[\alpha^{\mathsf{Min}}{=}0.10~,~\alpha^{\mathsf{Max}}{=}0.85]$ is a dynamic/instantaneous weighting coefficient.
 We model $\alpha^{\mathsf{Inst}}(t)$ as a simple function of the walker’s instantaneous accuracy $\text{A}^{\mathsf{Inst}}(t)$ as
@@
 As training progresses, memory influence gradually increases (green and blue regions), helping to mitigate model bias and yielding notable performance gains; especially in the middle to late stages of model training.
-3.4 Random Walks with Flexible Memory
+III-D Random Walks with Flexible Memory
 A key challenge for a random walker is forgetting patterns learned from previously visited nodes as it continues updating its model during new node visits, mirroring the catastrophic forgetting phenomenon in ML.
@@
 Future Research Directions The exploration of $\mathbb{X}$L with a single walker opens up various future directions, discussed below.
-Connection to Graph Signal Processing (GSP): We suggest exploring node importance metrics for $\mathbb{X}$L inspired by core signal processing methods, such as GSP [47, 48, 49, 50, 51].
+Connection to Graph Signal Processing (GSP): We suggest exploring node importance metrics for $\mathbb{X}$L inspired by core signal processing methods, such as GSP [13].
 GSP incorporates graph topology into node signal representation through eigen-decomposition of the graph Laplacian.
@@
 ML Convergence under Time-Varying Transition Matrices: Characterizing the ML convergence under time-varying transition matrices is crucial and will involve obtaining ML convergence bounds for convex and non-convex loss functions, considering the impacts of random walk trajectory on its model convergence.
-Comparison with Traditional Markov Chain Designs: Analyzing how the dynamic design of transition matrices differs from time-homogeneous Markov chains considering traditional metrics (e.g., covering time and mixing time [52, 53]) is interesting.
+Comparison with Traditional Markov Chain Designs: Analyzing how the dynamic design of transition matrices differs from time-homogeneous Markov chains considering traditional metrics (e.g., covering time and mixing time) is interesting.
 This involves demonstrating ‘how’ and ‘why’ traditional random walk visiting patterns may lead to inferior performance when compared to $\mathbb{X}$L-specific strategies.
@@
 The bottom plot presents the walkers’ final accuracies, highlighting an initial sharp gain followed by diminishing gains as more walkers are deployed.
-4 $\mathbb{X}$-Learning with Multiple Random Walkers
+IV $\mathbb{X}$-Learning with Multiple Random Walkers
 Expanding $\mathbb{X}$L, we consider scenarios involving multiple random walkers, opening up new research opportunities.
@@
 For instance, multiple walkers can have overlapping traversal patterns and potential collaborative or competitive dynamics as detailed next.
-4.1 Collaboration Among Walkers
+IV-A Collaboration Among Walkers
 Having multiple random walkers independently exploring the network can be inefficient, motivating us to study interactions between the walkers.
-Drawing from physics concepts on interacting particles [54, 55, 56, 57, 58] and epidemic spreading [59, 60, 61, 62, 63, 64], we explore cooperation between random walkers through information transfer.
+Drawing from physics concepts on interacting particles [14] and epidemic spreading [15], we explore cooperation between random walkers through information transfer.
 Specifically, when two walkers meet at a node, they “collide” and aggregate their models (e.g., via weighted averaging their models based on the number of seen datapoints).
@@
 Investigating the factors influencing this diminishing reward is an exciting research direction.
-Figure 6: Impact of inter-walker attraction on model performance.
-As the attraction between walkers increases (i.e., higher values of AA, leading to shorter inter-collision intervals), the walkers interact more frequently, which results in mitigating model bias and an improved performance.
-For comparison, the red curve illustrates the (hypothetical) upper bound achieved by centrally aggregating the walkers’ models after every jump.
-4.2 Attraction/Repulsion Between Multiple Random Walkers
+IV-B Attraction/Repulsion Between Multiple Random Walkers
 We next aim to control the frequency of collisions between walkers.
@@
 If a new walker arrives at a node where a trace is still fresh/reliable, it can aggregate its ML model with this trace, while accounting for the age of the trace.
-From Single-Task to Multi-Task Learning: The walkers can potentially train different ML tasks, introducing multi-task learning [65, 66, 67, 68, 69] into $\mathbb{X}$L.
+From Single-Task to Multi-Task Learning: The walkers can potentially train different ML tasks, introducing multi-task learning into $\mathbb{X}$L.
 This opens up exploration of (i) aggregation between the walkers based on their task similarity, (ii) optimal walker-to-task allocation, and (iii) resource optimization (e.g., number of SGD iterations of walkers).
@@
 Redefining the Notion of Centrality and Perception: Incorporating multiple random walkers necessitates reevaluating node centrality and walkers’ perceptions defined in Sec.
-3.
+III.
 These revamps should be done to avoid redundancy in exploration by assigning different centrality values to a node across walkers.
-5 Conclusion
+Figure 6: Impact of inter-walker attraction on model performance.
+As the attraction between walkers increases (i.e., higher values of AA, leading to shorter inter-collision intervals), the walkers interact more frequently, which results in mitigating model bias and an improved performance.
+For comparison, the red curve illustrates the (hypothetical) upper bound achieved by centrally aggregating the walkers’ models after every jump.
+V Conclusion
 We introduced $\mathbb{X}$L, a novel framework that expands distributed ML by treating ML models as autonomous agents capable of traversing the network through flexible, multi-hop strategies.
@@
 A series of future research directions were discussed, broadening the horizon of $\mathbb{X}$L.
-Acknowledgements
-Not applicable
-Funding
-S. Hosseinaliour was partially funded by NSF 2421761 while working on this paper.
-Competing interests
-The authors declare no competing interests.
-Additional information
-Correspondence and requests for materials should be addressed to S. Hosseinalipour.